OpenStudy (anonymous):

in a certain population of 540 cats, a single gene with two alleles controls thickness of the cats fur. allele C provided thick fur and is dominant, and c provides thin fur. in this population, 49 cats have thin fur. what are the frequencies of the two alleles?

OpenStudy (aaronq):

we find the frequency of homozygous recessive individuals \(\sf f(cc)=\dfrac{49}{540}=0.0907\approx 0.091\) From the expanded Hardy-weinberg equation we know that \(\sf c^2=f(cc)=0.091\rightarrow c=\sqrt{0.091}=0.30\) \(\sf f(C)=1-0.30=0.70\) Therefore, allele C is 0.7 and allele c is 0.3

OpenStudy (anonymous):

thank you. i have another question. would you mind helping me on it as well?

OpenStudy (aaronq):

no problem. sure post it

OpenStudy (anonymous):

hemophilia is it generally blood disorder in humans the gene responsible for its genes are linked recessive meaning it is present on the X chromosome when the gene is present in males XH it always expresses itself meaning that the most blood won't clot properly when present in females and a heterozygous concession condition XhXH the gene doesn't express itself but the female becomes a carrier meaning that her offspring could have hemophilia when presented in a homozygous recessive condition in females xhxh expresses itself one of every ten 10,000 people is affected by this disease what could be the allele frequency for this disease in the entire population @aaronq

OpenStudy (aaronq):

This one is more difficult, because the frequencies vary between males and females because of the allele requirement (only one in males). In females the homozygous recessive genotype \((X^hX^h)\) occurs in \(\dfrac{1}{10,000}\) \(\sf f(X^hX^h)=0.0001=(X^h)^2\rightarrow X^h=\sqrt{0.0001}=0.01 \) Try to figure out the rest, and remember the difference between males and female